Calculus I, Fall 2015 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 04

1. Section 3.5: problems 17, 49, 57, 58.

2. Section 3.6: Do as many as time permits from problems 51, 53, ⋯, 75.

3. Assume g(2) = 3, g′(2) = 0.1, f′(2) = 3, f′(3) = 4 and f′(4) = 5. What is \frac{d}{dx}f(g(x)) at x = 2?

4. Section 3.7: problems 27, 31, 42, 48.